OpenStudy (anonymous):

Find the slope of the line that passes through the pairs of points. (0, 0) (3,5) A.) 3/5 B.) 5/3 C.) -5/3 Please help ;(

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=5-0/3-0 m=5/3

OpenStudy (pottersheep):

^Yep, just remember slope is rise/run

OpenStudy (anonymous):

@Mariam.A thank you